OpenStudy (anonymous):

Consider the functions f(x) = x2 − 13 and g(x) = x + 5. What is the value of (f + g)(3)?

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

ok you can add both functions first and then substitute the x for 3 or you can substitute on both functions the x for 3 and then add the results

OpenStudy (anonymous):

Question is quite simple since both are functions of x then (f+g)(x)=f(x)+g(x) so (f+g)(x)=2x-13+x+5 (f+g)(x)=3x-8 (f+g)(3)=3x3-8=9-8=1

OpenStudy (ivettef365):

ipm sorry but that is wrong because is no 2x, it is x^2

OpenStudy (anonymous):

ya 1 is not an answer

OpenStudy (ivettef365):

so you have x^2 -13 +x +5

OpenStudy (ivettef365):

that is equal to x^2 +x -8 now substitute the x for 3 and what do you get

OpenStudy (anonymous):

@ivettef365 not my fault in the question it was not given raised to the power

OpenStudy (anonymous):

sorry

OpenStudy (ivettef365):

so what are the equations

OpenStudy (anonymous):

-32?

OpenStudy (anonymous):

its x^2

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

you are talking about f(x) right

OpenStudy (anonymous):

no the x right after that

OpenStudy (ivettef365):

ok the problems are: f(x) = x^2 -13 and g(x) = x +5

OpenStudy (anonymous):

yes

OpenStudy (ivettef365):

ok, then it was correct all alone, so we have x^2 +x-8, substitute x for 3 what do you get

OpenStudy (anonymous):

-4?

OpenStudy (ivettef365):

actually positive 4

OpenStudy (ivettef365):

you have (3)^2 +(3) - 8 9 +3 -8 = ??

OpenStudy (anonymous):

4 becuz the 12 is bigger

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

yes! ty

OpenStudy (ivettef365):

yw :)